OpenStudy (anonymous):

Use la grange multipliers to find the indicated extrema, assuming that x and y are positive. f(x,y) = x^2-y^2 constraint = g(x,y): 2y-x^2=0

OpenStudy (anonymous):

Partial derivatives of f and g \[\Delta f(x,y)= <2x,-2y>\] \[\lambda \Delta g(x,y)=<-2\lambda x,2\lambda>\]

OpenStudy (anonymous):

\[2x=-2 lambda x\] \[-2y=2\lambda\]

OpenStudy (anonymous):

x would be \[\lambda x\] and y \[2\lambda\] right?

OpenStudy (anonymous):

oh wait I mean y would be just lambda

OpenStudy (anonymous):

\[\nabla f=\lambda \nabla g\]\[\left(\begin{matrix}2x \\ -2y\end{matrix}\right)=\lambda \left(\begin{matrix}-2x \\ 2\end{matrix}\right)\]\[2x=-2x \lambda\]simplifying gives lambda=-1\[-2y=2 \lambda\] simplifying gives y=-lambda, y=1 from your constraint\[2y-x ^{2}=0\]\[x ^{2}=2\]\[x=\pm \sqrt{2}\]

OpenStudy (anonymous):

Oh okay I get it now Thank you.